OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or equal to x - 5 y less than or equal to -x - 4

OpenStudy (anonymous):

A. (1, 10) B. (-1, 10) C. (10, 1) D. (1, -10)

OpenStudy (anonymous):

y<=x-5 y<=-x-4 2y<=-9 y<=-4/5 then y=-10 is valid and x = -1 too for it so D.

OpenStudy (anonymous):

thank you